# PII-Scope: A Benchmark for Training Data PII Leakage Assessment in LLMs

### v1

@@ Abstract
In this work, we introduce PII-Scope, a comprehensive benchmark designed to evaluate state-of-the-art methodologies for PII extraction attacks targeting LLMs across diverse threat settings. Our study provides a deeper understanding of these attacks by uncovering several hyperparameters (e.g., demonstration selection) crucial to their effectiveness. Building on this understanding, we extend our study to more realistic attack scenarios, exploring PII attacks that employ advanced adversarial strategies, including repeated and diverse querying, and leveraging iterative learning for continual PII extraction. Through extensive experimentation, our results reveal a notable underestimation of PII leakage in existing single-query attacks. In fact, we show that with sophisticated adversarial capabilities and a limited query budget, PII extraction rates can increase by up to fivefold when targeting the pretrained model. Moreover, we evaluate PII leakage on finetuned models, showing that they are more vulnerable to leakage than pretrained models. Overall, our work establishes a rigorous empirical benchmark for PII extraction attacks in realistic threat scenarios and provides a strong foundation for developing effective mitigation strategies.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated a tendency to memorize training data, which includes not only benign useful content but also sensitive Personal Identifiable Information (PII) of individuals that are part of the pretraining dataset. Notably, LLMs are typically pretrained on a vast corpus of data scraped from the internet, which can contain sensitive PII. With the recent enforcement of regulations such as the AI Act (European Commission, 2021) and GDPR (Parliament & of the European Union, 2016), ensuring the privacy of data subjects has become paramount.
Due to growing privacy concerns, early research (Carlini et al., 2021a; 2022) primarily focused on the memorization of general, non-sensitive suffixes, while more recent studies (Lukas et al., 2023; Nakka et al., 2024; Kim et al., 2024; Huang et al., 2022a) have specifically investigated the memorization of PII, highlighting the significant privacy risks associated with this phenomenon. However, these studies often vary in their experimental setups and assumptions regarding data access, leading to unstandardized comparisons across studies. Consequently, a clear and unified understanding of PII extraction attacks has not been established yet in the literature. Furthermore, several works (Sun et al., 2024; Wang et al., 2023) have evaluated privacy leakage as part of the larger goal of assessing LLM trustworthiness including safety, harmfulness, and other hazards (Vidgen et al., 2024). These studies, however, are limited to few isolated privacy attack scenarios from Huang et al. (Huang et al., 2022a), highlighting a crucial absence of comprehensive evaluations. This situation underscores the urgent need for critical benchmarking of PII attacks to effectively assess and mitigate PII leakage.
To address these critical gaps, we present PII-Scope, the first comprehensive empirical assessment of PII extraction attacks from pretrained LLMs. First, we conduct a systematic analysis of potential PII attacks within each threat scenario and examine the sensitivity of the corresponding attack methodologies. Building on these insights, we further explore PII attacks using advanced attacking capabilities. Our key contributions are as follows: 1) We propose a taxonomy of PII attacks, categorizing them based on the threat model and data accessibility assumptions. 2) We provide an in-depth analysis of each attack’s sensitivity to its internal attack hyperparameters. 3) We develop PII-Scope, a realistic and standardized evaluation methodology of these attacks. 4) Finally, PII-Scope demonstrates that current PII attack approaches significantly underestimate PII leakage and shows that extraction rates can improve by up to threefold with a limited query budget.
@@ 3 PII Attacks Taxonomy
To enable a detailed analysis of PII attacks, we categorize current PII attacks in the literature based on two key dimensions: access to the model and access to the pretraining dataset. Figure 1 illustrates the categorization of threat settings and the potential PII attacks within each setting. We distinguish between black-box and white-box settings (i.e., whether the attacker has access to the target LLM’s parameters) at the first level, and consider the attacker’s access to the pretraining data at the second level. The latter can occur at three distinct levels: 1. access to the true prefix of the query data subject, 2. knowledge of PII pairs related to a few other data subjects included in the pretraining dataset, and 3. access to the true prefixes of a few other data subjects in the pretraining dataset different from the target data subject.
Figure 2: Illustration of input prompt construction with different PII attacks. The attacker employs various strategies, including: prompting the model with true prefixes (Carlini et al., 2021a); using template prompts (Huang et al., 2022a); leveraging additional context from PII pairs (ICL) (Huang et al., 2022a), true prefixes of other data subjects (PII Compass) (Nakka et al., 2024); or learning soft prompt on a small subset containing PII pairs of a few data subjects (Kim et al., 2024).
@@
True-prefix $r_{q}$ True-prefix of a subject $q$ in Evaluation PII dataset
@@
True-prefix $r_{j}^{*}$ True-prefix of a subject $j$ in Adversary PII dataset
Subject name $s_{j}^{*}$ Name of a subject $j$ in Adversary PII dataset
Subject PII $p_{j}^{*}$ Email PII of a subject $j$ in Adversary PII dataset
@@
    Remark. Additionally, we note that the SPT attack, originally proposed in (Kim et al., 2024), shares similarities with the recent SOS attack (Yang et al., 2024), though applied in a slightly different context. In the SOS attack (Yang et al., 2024), the authors demonstrate that soft prompt embeddings, finetuned on as few as 10 examples, can jailbreak aligned LLMs to generate unsafe content. This highlights the versatility of finetuning-based attacks, not only in PII extraction but also in broader safety-related attacks. Moreover, recent studies (Lermen et al., 2023; Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023; Halawi et al., 2024) have explored malicious model finetuning to break LLM alignment. Overall, attacks based on finetuning, whether through parameter-efficient methods like soft-prompt tuning (Lester et al., 2021) or direct finetuning on a handful of examples, represent an emerging paradigm applicable to a wide range of tasks for jailbreaking LLM alignment, and our current focus SPT attack (Kim et al., 2024) can be viewed as one example attack in this family of finetuning-based attacks.
@@
". Obviously that is counter to your knowledge of the transaction.
@@
Figure 5: Existing benchmark. Among the 3,333 data subjects in the original Enron PII leakage dataset (Huang et al., 2022a), there are only 404 unique email domains, indicating that many data subjects share domains. Here, we show the frequency of the top-30 occurring email domains out of the 404, along with the cumulative proportion these data subjects constitute in the original dataset. We observe that just the top-30 domains account for 45% of the data subjects in the original dataset.
@@ 4 Experimental Setting
Following the overview of the PII attacks, we now turn our attention to the benchmark assessment set, which is crucial for evaluating these attacks.
Benchmark Dataset. The original Enron PII leakage assessment dataset (Huang et al., 2022a) contains 3,333 non-Enron data subjects, each with a name and email pair. Upon exploring this dataset, we observed significant email-domain overlap among the data subjects. Despite the dataset comprising 3,333 data points, there were only 404 unique email domains. Figure 5 illustrates the frequency of the top-30 email domains out of 404 domains, which account for almost 45% of the data subjects. Additionally, the user-part of the email PII is often confined to a few predictable patterns, meaning that knowing the domain-part can make extracting the full email PII much easier, almost a trivial task.
@@
To address these concerns, we curated a pruned dataset comprising 404 data subjects, each uniquely associated with a specific domain (404 domains in total). After manual inspection, we excluded 32 data subjects due to either short or unclear single-word names (eg., subject names such as "s", "Chris", "Sonia"). The remaining 372 data subjects were then divided into two groups: $M=64$ subjects designated for attacker access (used in ICL or SPT attacks) are grouped under $\mathcal{D}_{\text{adv}}$, and the remaining $N=308$ subjects, intended for unbiased evaluation, are grouped under $\mathcal{D}_{\text{eval}}$.
Model. All experiments are conducted on single GPT-J-6B (Wang & Komatsuzaki, 2021), a standard model for evaluating PII leakage, chosen due to the publicly available information about its pretraining dataset. For reproducibility, we provide detailed information about the 372 data subjects used for our experiments, along with further implementation details of each PII attack in Appendix A.1.
@@ 5 Sensitivity of PII Attacks
From this section, we shift our focus to the empirical evaluation of PII attacks. To critically understand the strengths and weaknesses of each attack, we first systematically investigate the robustness of each PII attack with regard to its internal hyperparameters in single-query budget, i.e., LLM is queried only once per query data subject.
@@
Figure 6: Performance of the true-prefix attack on the pretrained model. The PII extraction rate of the true-prefix attack (Carlini et al., 2021a) increases with the number of tokens in the true prefix, with performance starting to saturate after 100 tokens.
@@
Figure 7: Sensitivity of hard-prompt attacks on the pretrained model. (a) The template attack (Huang et al., 2022a) shows sensitivity to the prompt template structure, (b) the ICL attack (Huang et al., 2022a) demonstrates sensitivity to the selection of demonstrations (observable by the large confidence intervals), and (c) the PII Compass attack (Nakka et al., 2024) reveals the impact of varying context sizes with true prefixes from $\mathcal{D}_{{adv}}$.
@@ 5.4 PII Compass Attack
Figure 7(c) shows the extraction rates across the 64 different choices of $r_{j}^{*}$, further stratified by different prefix lengths $L=\{25,50,100\}$. We observe significant variance in extraction rates, with differences as large as 8% as $r_{j}^{*}$ varies. This suggests that extraction performance highly depends on the specific $r_{j}^{*}$ used. A well-chosen $r_{j}^{*}$ can yield extraction rates as high as 8%, while a poor choice may result in performance even lower than the baseline template attack using $T_{q}$ alone, as shown in Figure 7(a). Each vertical boxplot in Figure 7(c) represents the distribution of extraction rates obtained using $M=64$ different true-prefixes $\{r_{j}^{*}\}_{j=1}^{M=64}$ for a given prefix length.
Interestingly, the number of tokens in the true-prefix $r_{j}^{*}$ has minimal impact on performance. Even with $L=25$ tokens, sufficient contextual information exists to ground the victim LLM $f$ effectively, achieving performance similar to that of larger token lengths, such as $L=150$.
@@
Figure 8: Sensitivity of SPT Attack (Kim et al., 2024) on the pretrained model. We examine the variation in PII extraction rates by analyzing the impact of three independent factors. Each factor is varied independently from the base configuration, and the results show that the SPT attack requires careful hyperparameter selection for optimal performance.
@@ 5.5 Soft-Prompt Tuning Attacks
Training soft prompt embeddings in the SPT attack involves multiple hyperparameters, such as the number of tokens in the soft prompt, initialization method, and the number of training epochs. To better isolate the impact of each, we vary these hyperparameters independently from the base configuration. For the base configuration, we use a task-aware prompt initialization string: ‘‘Extract the email address associated with the given name’’, with the number of tokens in the soft prompt $L$ set to 50 and the number of training epochs set to 20 (see Appendix A.2 for more details).
Impact of Number of Tokens in the Soft Prompt. We vary the number of tokens of the soft prompt $L$ from 20 to 120. The results, shown in Figure 8(a), indicate that performance improves as the number of tokens in the soft prompt increases, peaking between 40 and 60 tokens, after which performance declines sharply. This degradation with larger numbers of tokens in the soft prompt is attributed to overfitting on the small set of data subjects in $\mathcal{D}_{\text{adv}}$ used for training.
@@ 4 Templates: [A, B, C, D])
Table 4: Evaluating PII attacks with higher query budgets on the pretrained model. The first four columns outline the threat setting in terms of data access in $\mathcal{D}_{adv}$ and $\mathcal{D}_{eval}$. The fifth column shows the model access type (W.B.: white box, B.B.: black box). We conduct PII attacks by querying the model multiple times, either through simple top-$k$ model sampling or by varying configuration settings within each attack method. Overall, we observe that the extraction rate improves by approximately 1.37x - 5.38x compared to the best extraction rate observed with a single query.
@@
Despite the significantly increased extraction rates across all methods, it is crucial to emphasize that each attack involves several sensitive hyperparameters, as discussed in §$\S$5. Therefore, making direct comparisons between PII attack methods at a fixed query budget may introduce bias due to confounding factors. Nevertheless, the primary goal of this experiment is to demonstrate that, in real-world scenarios, an adversary could leverage these insights to substantially enhance PII extraction rates of at least once in K queries—by 1.3x - 5.4x times compared to the best rates achieved in a single-query setting. It is important to note that the predictions generated with $K$ queries represent only the candidate PIIs of the query data subject, which may include the ground-truth PII. The attacker would need to perform additional work to identify the actual ground-truth PII among these $K$ predictions. This could be achieved either by applying ranking metrics (eg., loss (Yeom et al., 2018), Zlib (Carlini et al., 2021b)) or through manual verification.
Figure 10: Continual PII Extraction on the pretrained model. We report the extraction rates of the SPT attack Kim et al. (2024) over ten rounds for four templates in a continual learning setting. At the end of each round, successfully extracted PIIs are incorporated to retrain the soft prompt embeddings for the subsequent round. The average extraction rate, along with its range, is plotted for the first five soft-prompt initializations shown in Figure 19.

### v2

# PII-Scope: A Comprehensive Study on Training Data PII Extraction Attacks in LLMs

@@ Abstract
In this work, we present PII-Scope, a comprehensive study benchmarking state-of-the-art methodologies for personally identifiable information (PII) extraction attacks targeting LLMs across diverse threat settings. Our study provides a deeper understanding of these attacks by uncovering several hyperparameters (e.g., demonstration selection) crucial to their effectiveness. Building on this understanding, we extend our study to more realistic attack scenarios, exploring PII attacks that employ advanced adversarial strategies, including repeated and diverse querying, and leveraging iterative learning for continual PII extraction. Through extensive experimentation, our results reveal a notable underestimation of PII leakage in existing single-query attacks. In fact, we show that with sophisticated adversarial capabilities and a limited query budget, PII extraction rates can increase by up to fivefold when targeting the pretrained model. Moreover, we evaluate PII leakage on finetuned models, showing that they are more vulnerable to leakage than pretrained models. Overall, our work establishes a rigorous empirical benchmark for PII extraction attacks in realistic threat scenarios and provides a strong foundation for developing effective mitigation strategies.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated a tendency to memorize training data, which ranges from benign and valuable knowledge to unintentionally embedded personal information. Notably, since LLMs are usually pretrained on vast datasets collected from the internet, which inevitably contain sensitive personally identifiable information (PII), there is a risk that the models memorize and unintentionally reveal this information during inference. With the recent enforcement of regulations such as the AI Act (European Commission, 2021) and GDPR (Parliament & of the European Union, 2016), ensuring the privacy of data subjects has become paramount.
Due to growing privacy concerns, early research (Carlini et al., 2021a; 2022) primarily focused on the memorization of general, non-sensitive suffixes, while more recent studies (Lukas et al., 2023; Nakka et al., 2024; Kim et al., 2024; Huang et al., 2022a) have specifically investigated the memorization of PIIs, highlighting the significant privacy risks associated with this phenomenon. However, these studies often vary in their experimental setups and assumptions regarding the threat model and data access, leading to unstandardized comparisons across studies. At present, the literature has not yet reached a clear and unified understanding of PII extraction attacks. Furthermore, while several works (Sun et al., 2024; Wang et al., 2023) have evaluated privacy leakage as part of the larger goal of assessing LLM trustworthiness including safety, harmfulness, and other hazards (Vidgen et al., 2024), these studies are limited to few isolated privacy attack scenarios from Huang et al. (Huang et al., 2022a), highlighting a crucial absence of comprehensive evaluations. To summarize, current situations underscore the urgent need for critical benchmarking of PII attacks to effectively assess and mitigate PII leakage.
To address these critical gaps, we present PII-Scope, a first comprehensive study on the empirical assessment of PII extraction attacks from LLMs. First, we conduct a systematic analysis of potential PII attacks within each threat scenario and examine the sensitivity of the corresponding attack methodologies. Building on these insights, we further explore PII attacks using advanced attacking capabilities. Our key contributions are as follows:
  1. 1.
    We propose a taxonomy of PII attacks, categorizing them based on the threat model and data accessibility assumptions.
  2. 2.
    We provide an in-depth analysis of each attack’s sensitivity to its internal attack hyperparameters.
  3. 3.
    We provide a realistic and standardized evaluation methodology of PII attacks and demonstrates that current PII attack approaches significantly underestimate PII leakage in single-query settings and shows that extraction rates can improve by up to fivefold in multi-query attack settings with adversarial strategies.
  4. 4.
    We conduct extensive experiments on two LLMs, GPT-J 6B (Wang & Komatsuzaki, 2021) and Pythia 6.9B (Biderman et al., 2023), against two PIIs, email and phone numbers, in both pretrained and finetuned settings. We validate the higher extraction rates with advanced attacker capabilities in multi-query attack settings.
@@ 2 Related Work
3 Overview of PII-Scope
To comprehensively assess the strengths and limitations of PII extraction attacks, we introduce PII-Scope, a standardized evaluation framework designed to examine these attacks. By using this framework, we explore the intricate interplay of PII leakage rates across different threat settings, investigating how leakage varies in diverse contexts. Our experiments are conducted from two perspectives: the attack perspective, to understand the factors influencing attack success, and the model perspective, to assess leakage rates under advanced attacker capabilities in higher-query attack settings.
We begin in Section 4 by presenting the taxonomy of PII attacks and discussing the details of each attack. In Section 5, we describe our benchmark dataset used in the experiments, focusing on preventing data contamination when evaluating PII attacks. Next, in Section 6, we demonstrate that current single-query attacks are sensitive to hyperparameter design choices within the attack. To better understand the extent of PII leakage, Section 7 explores leakage rates in higher query settings and a novel continual learning setting. In Section 8, we provide ablation studies to further analyze the factors affecting PII leakage rates. To generalize our findings, Section 10 presents results using additional LLMs, and Section 11 evaluates additional PIIs. Finally, we also show that the underestimation of leakage generalizes when the models are finetuned with PII-scrubbing, a popular defense in the pretraining scenario.
Through our extensive experiments, we show the following key findings: 1. Single-query attacks are highly sensitive to design choices (see Section  6). 2. Multi-query attacks show increased leakage rates, up to five-fold (see Sections 7 and 10). 3. Finetuned models are more vulnerable to PII attacks (see Sections 910 and  11).
@@ 4 PII Attacks Taxonomy
To enable a detailed analysis of PII attacks, we categorize current PII attacks in the literature based on two key dimensions: access to the model and access to the pretraining dataset. Figure 1 illustrates the categorization of threat settings and the potential PII attacks within each setting. We distinguish between black-box and white-box settings (i.e., whether the attacker has access to the target LLM’s parameters) at the first level, and consider the attacker’s access to the pretraining data at the second level. The latter can occur at three distinct levels: 1. access to the true training data prefix of the query data subject, 2. knowledge of PII pairs related to a few other data subjects included in the pretraining dataset, and 3. access to the true training data prefixes of a few other data subjects that are different from the target data subject.
Figure 2: Illustration of input prompt construction with different PII attacks. The attacker employs various strategies, including prompting the model with true prefixes (Carlini et al., 2021a); using template prompts (Huang et al., 2022a); leveraging additional context from PII pairs (ICL) (Huang et al., 2022a), true prefixes of other data subjects (PII Compass) (Nakka et al., 2024); or learning soft prompt on a small subset containing PII pairs of a few data subjects (Kim et al., 2024).
@@
True-prefix $r_{j}^{*}$ True prefix of a subject $j$ in Adversary PII dataset
Subject name $s_{j}^{*}$ Name of a subject $j$ in Adversary PII dataset
Subject PII $p_{j}^{*}$ Email PII of a subject $j$ in Adversary PII dataset
True-prefix $r_{q}$ True prefix of a subject $q$ in Evaluation PII dataset
@@ 4.5 SPT Attack
". Obviously, that is counter to your knowledge of the transaction.
@@
Figure 5: Existing benchmark. Among the 3,333 data subjects in the original Enron PII leakage dataset (Huang et al., 2022a), there are only 404 unique email domains, indicating that many data subjects share the same domains. Here, we show the frequency of the top-30 most common email domains from the 404 domains, along with the cumulative proportion these data subjects constitute in the original dataset. We observe that just the top-30 domains alone account for 45% of the data subjects in the original dataset.
@@ 5 Experimental Setting
Following the overview of the PII attacks, we now turn our attention to the benchmark assessment set, which is crucial for evaluating these attacks. We conduct our experiments to extract two PII, emails and Phone numbers.
Email PII Benchmark Dataset. The original Enron PII leakage assessment dataset (Huang et al., 2022a) contains 3,333 non-Enron data subjects, each with a name and email pair. Upon exploring this dataset, we observed significant email-domain overlap among the data subjects. Despite the dataset comprising 3,333 data points, there were only 404 unique email domains. Figure 5 illustrates the frequency of the top-30 email domains out of 404 domains, which account for almost 45% of the data subjects. Additionally, the user-part of the email PII is often confined to a few predictable patterns, meaning that knowing the domain-part can make extracting the full email PII much easier, almost a trivial task.
@@
To address these concerns, we curated a pruned dataset comprising 404 data subjects, each uniquely associated with a specific domain (404 domains in total). After manual inspection, we excluded 32 data subjects due to either short or unclear single-word names (eg., subject names such as "s", "Chris", "Sonia"). The remaining 372 data subjects were then divided into two groups: $M=64$ subjects designated for attacker access (used in ICL or SPT attacks) are grouped under $\mathcal{D}_{\text{adv}}$, and the remaining $N=308$ subjects, intended for unbiased evaluation, are grouped under $\mathcal{D}_{\text{eval}}$. For reproducibility, we provide detailed information about the 372 data subjects used for our experiments, along with further implementation details of each PII attack in Appendix A.1.
Phone Number PII Benchmark Dataset. We randomly selected 500 data subjects from 2078 subjects in the evaluation set provided by the authors in the ICL attack (Shao et al., 2023). We then set aside 64 data subjects to form the adversary dataset $D_{adv}$ and evaluate the extraction rates on the remaining 436 data subjects.
Models. We run our experiments on three LLMs. All the main experiments are conducted on GPT-J-6B (Wang & Komatsuzaki, 2021), a standard model for evaluating PII leakage, chosen due to the publicly available information about its pretraining dataset, the PILE corpus. We also evaluate attacks on Pythia 6.9B (Biderman et al., 2023), which is also pretrained on PILE. Finally, to show the generalization of our findings, we evaluate attacks on the LLaMa 7B model (Touvron et al., 2023), finetuned on the Enron email dataset.
@@ 6 Sensitivity of PII Attacks
In this section, we shift our focus to the empirical evaluation of PII attacks. As a representative study, we first focus on email PII extraction evaluations on GPT-J 6B (Wang & Komatsuzaki, 2021), providing an in-depth analysis. We then expand our evaluations to Pythia 6.9B (Biderman et al., 2023) in Section 10, and to phone number PII in Section 11.
To first critically understand the strengths and weaknesses of each attack, we first systematically investigate the robustness of each PII attack with regard to its internal hyperparameters in single-query budget, i.e., LLM is queried only once per query data subject.
@@
Figure 6: Performance of the true-prefix Email PII attack on the pretrained GPTJ-6B model. The PII extraction rate of the true-prefix attack (Carlini et al., 2021a) increases with the number of tokens in the true prefix, with performance starting to saturate after 100 tokens.
@@
Figure 7: Sensitivity of hard-prompt Email PII attacks on the pretrained GPTJ-6B model. (a) The template attack (Huang et al., 2022a) shows sensitivity to the prompt template structure, (b) the ICL attack (Huang et al., 2022a) demonstrates sensitivity to the selection of demonstrations (observable by the large confidence intervals), and (c) the PII Compass attack (Nakka et al., 2024) reveals the impact of varying context sizes with true prefixes from $\mathcal{D}_{{adv}}$.
@@ 6.4 PII Compass Attack
Figure 7(c) shows the extraction rates across the 64 different choices of $r_{j}^{*}$, further stratified by different prefix lengths $L=\{5,10,25,50,100\}$. We observe significant variance in extraction rates, with differences as large as 8% as $r_{j}^{*}$ varies. This suggests that extraction performance highly depends on the specific $r_{j}^{*}$ used. A well-chosen $r_{j}^{*}$ can yield extraction rates as high as 8%, while a poor choice may result in performance even lower than the baseline template attack using $T_{q}$ alone, as shown in Figure 7(a). Each vertical boxplot in Figure 7(c) represents the distribution of extraction rates obtained using $M=64$ different true-prefixes $\{r_{j}^{*}\}_{j=1}^{M=64}$ for a given prefix length.
[…]
@@
Figure 8: Sensitivity of SPT Attack (Kim et al., 2024) on the pretrained GPTJ-6B model for email PII extraction. We examine the variation in PII extraction rates by analyzing the impact of three independent factors. Each factor is varied independently from the base configuration, and the results show that the SPT attack requires careful hyperparameter selection for optimal performance.
@@ 6.5 Soft-Prompt Tuning Attacks
Training soft prompt embeddings in the SPT attack involves multiple hyperparameters, such as the number of tokens in the soft prompt, the initialization method, and the number of training epochs. To better isolate the impact of each, we vary these hyperparameters independently from the base configuration. For the base configuration, we use a task-aware prompt initialization string: ‘‘Extract the email address associated with the given name’’, with the number of tokens in the soft prompt $L$ set to 50 and the number of training epochs set to 20 (see Appendix A.2 for more details).
Impact of Number of Tokens in the Soft Prompt. We vary the number of tokens of the soft prompt $L$ from 20 to 120. The results, shown in Figure 8(a), indicate that performance improves as the number of tokens in the soft prompt increases, peaking between 40 and 60 tokens, after which performance declines.
@@ 4 Templates: [A, B, C, D])
Table 4: Evaluating Email PII attacks with higher query budgets on the pretrained GPTJ-6B (Wang & Komatsuzaki, 2021). The first four columns outline the threat setting in terms of data access in $\mathcal{D}_{adv}$ and $\mathcal{D}_{eval}$. The fifth column shows the model access type (W.B.: white box, B.B.: black box). We conduct PII attacks by querying the model multiple times, either through simple top-$k$ model sampling or by varying configuration settings within each attack method. Overall, we observe that the extraction rate improves by approximately 1.37x - 5.38x compared to the best extraction rate observed with a single query.
@@
Despite the significantly increased extraction rates across all methods, it is crucial to emphasize that each attack involves several sensitive hyperparameters, as discussed in §$\S$6. Therefore, making direct comparisons between PII attack methods at a fixed query budget may introduce bias due to confounding factors. Nevertheless, the primary goal of this experiment is to demonstrate that, in real-world scenarios, an adversary could leverage these insights to substantially enhance PII extraction rates of at least once in K queries—by 1.3x - 5.4x times compared to the best rates achieved in a single-query setting. It is important to note that the predictions generated with $K$ queries represent only the candidate PIIs of the query data subject, which may include the ground-truth PII. The attacker would need to perform additional work to identify the actual ground-truth PII among these $K$ predictions. This could be achieved either by applying ranking metrics (eg., loss (Yeom et al., 2018), Zlib (Carlini et al., 2021b)) or through manual verification.
Figure 10: Continual Email PII Extraction on the pretrained GPTJ-6B (Wang & Komatsuzaki, 2021). We report the extraction rates of the SPT attack Kim et al. (2024) over ten rounds for four templates in a continual learning setting. At the end of each round, successfully extracted PIIs are incorporated to retrain the soft prompt embeddings for the subsequent round. The average extraction rate, along with its range, is plotted for the first five soft-prompt initializations shown in Figure 22.